find and equation of the line containing the given pair of points
(2,3) and (6,5)
y=?

Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

find and equation of the line containing the given pair of points
(2,3) and (6,5)
-------------------
slope = (5-3)/(6-2) = 2/4 = 1/2
-------
Linear Equation Form:
y = mx + b
Substitute and solve for "b":
3 = (1/2)2 + b
3 = 1 + b
b = 2
----
Equation:
y = (1/2)x+2
==================
cheers,
Stan H.

Answer by rfer(16322)